OpenStudy (anonymous):

What is the probability of getting no more than 2 wrong by guessing on a 8 question True-False test?

OpenStudy (anonymous):

So on an 8 question test with each question having two possible answers(true/false) you have 2x2x2x2x2x2x2x2 or 2^8 ways of answering the test. To get no more than 2 wrong means that you can get 0 wrong, 1 wrong, or 2 wrong. So from the total number of possible ways you need to subtract the ways that have more than 2 wrong and divide that number by the total number of possible ways. Does that help?

OpenStudy (anonymous):

so would it be something like (8C0 +8C1 +8C2)/(2^8)?

OpenStudy (anonymous):

Hmm. It's been so long. Let me dwell on this for a bit. Don't necessarily take my advice lol

OpenStudy (anonymous):

Have you done binomial distributions? http://www.mathwords.com/b/binomial_probability_formula.htm

OpenStudy (anonymous):

yes we have

OpenStudy (callisto):

Dunno if it works... No more than 2 => 0/ 1/ 2 wrong. For none of the choices are wrong, P = 8C0 (0.5)^8 (0.5)^(8-8) = 0.5^8 For 1 of the choices is wrong, P = 8C1 (0.5)^7 (0.5)^(8-7) = 8 ( 0.5)^8 For 2 of the choices are wrong, P = 8C2 (0.5)^6 (0.5) ^ (8-6) = 28 ( 0.5)^8 Probability required = ( 0.5)^8 + 8 ( 0.5)^8 + 28 ( 0.5)^8 ...... (hmm... my first trial using this method... DON'T trust me !!!)

OpenStudy (anonymous):

that's how its done in the answer! but can you explain how you did for the none wrong?

OpenStudy (callisto):

None is wrong => n=8, r=0 <-- None is wrong.. Then, 50% for correct, since it's all correct => 0.5 ^8 50% for incorrect, since None are incorrect => 0.5^(8-8) = 0.5^0 =1 Multiply all of them 8C0 x 0.5^8 x 1 = 0.5^8

OpenStudy (anonymous):

Thanks. That helps me too :)

OpenStudy (callisto):

Welcome!~~~

OpenStudy (anonymous):

thanks! you just saved my life!